Need help!!!!!
Solve for x. 3x – 2y = 6

Found 3 solutions by checkley71, stanbon, Nate:
Answer by checkley71(8403)

3X-2Y=6 OR 3X=2Y+6 OR X=(2Y+6)/3

Answer by stanbon(75887)

3x – 2y = 6
You would need a 2nd equation in x and y to arrive
at a unique solution.
As it is; no unique solution.
Cheers,
stan H.

Answer by Nate(3500)

3x – 2y = 6 original
3x - 2y + 2y = 6 + 2y add 2y to each side
3x = 6 + 2y combine
(3x)/3 = (6 + 2y)/3 divide each side by 3
x = 2 + (2/3)y